# UniGAP: A Universal and Adaptive Graph Upsampling Approach to Mitigate Over-Smoothing in Node Classification Tasks

### v1

@@ 4.2.1. Overall Performance Comparison (RQ1)
In summary, compared to all the baselines and datasets, UniGAP demonstrates a significant improvement, with an average absolute boost ranging from 1.4% to 6.7%, surpassing the SOTA baselines, showcasing the effectiveness of UniGAP.
@@
Homophily 0.66 0.63 0.77 0.42 0.00 0.02 0.05 0.00 Avg
GT (Dwivedi and Bresson, 2020) 79.08 ± 0.4 70.10 ± 0.8 82.22 ± 0.6 70.63 ± 0.4 84.18 ± 5.4 80.16 ± 5.1 82.74 ± 6.0 34.28 ± 0.7 72.92
GraphGPS (Rampášek et al., 2023) 79.94 ± 0.3 72.40 ± 0.6 82.44 ± 0.6 70.97 ± 0.4 82.21 ± 6.9 82.06 ± 5.1 85.36 ± 4.2 36.01 ± 0.9 73.92
NAGphormer (Chen et al., 2023) 80.57 ± 0.3 72.80 ± 0.8 84.20 ± 0.5 70.13 ± 0.6 80.12 ± 5.5 79.89 ± 7.1 82.97 ± 3.2 34.24 ± 0.9 73.11
@@
+HalfHop 77.42 ± 0.6 72.40 ± 0.7 81.36 ± 0.6 72.59 ± 0.2 71.62 ± 3.9 72.39 ± 4.8 66.14 ± 6.6 30.69 ± 0.7 68.07
+UniGAP 80.06 ± 0.6 73.10 ± 0.6 84.20 ± 0.5 72.97 ± 0.2 77.37 ± 3.5 75.95 ± 4.7 69.59 ± 6.2 33.73 ± 0.9 70.87
GAT (Velickovic et al., 2018) 77.81 ± 0.4 71.00 ± 0.8 83.18 ± 0.5 71.95 ± 0.4 60.46 ± 6.2 58.22 ± 3.7 63.59 ± 6.1 30.36 ± 0.9 64.57
+HalfHop 78.03 ± 0.5 71.80 ± 0.8 82.29 ± 0.8 72.56 ± 0.4 74.54 ± 6.1 66.48 ± 5.5 70.17 ± 6.0 30.28 ± 1.0 68.27
+UniGAP 80.10 ± 0.3 73.20 ± 0.5 84.23 ± 0.6 74.06 ± 0.3 78.64 ± 4.9 76.04 ± 5.1 74.81 ± 5.3 32.79 ± 0.8 71.73
GraphSAGE (Hamilton et al., 2017) 78.67 ± 0.4 71.80 ± 0.6 83.76 ± 0.5 71.49 ± 0.3 82.43 ± 6.1 75.95 ± 5.3 81.18 ± 5.5 34.23 ± 1.0 72.44
+HalfHop 78.94 ± 0.3 72.00 ± 0.8 84.27 ± 0.6 72.93 ± 0.5 85.95 ± 6.4 74.60 ± 6.0 85.88 ± 4.0 36.82 ± 0.7 73.92
+UniGAP 80.21 ± 0.7 73.60 ± 0.3 86.46 ± 0.4 73.71 ± 0.2 86.52 ± 4.8 83.21 ± 5.7 86.98 ± 3.2 37.14 ± 0.8 75.98
@@
+HalfHop 78.96 ± 0.9 72.40 ± 0.7 84.91 ± 0.6 72.91 ± 0.3 79.84 ± 6.1 80.04 ± 5.1 82.96 ± 3.7 36.01 ± 0.9 73.50
+UniGAP 80.79 ± 0.7 73.70 ± 0.4 86.22 ± 0.4 73.32 ± 0.3 82.08 ± 4.2 81.16 ± 5.0 84.06 ± 3.8 36.47 ± 1.1 74.76
@@
+UniGAP 80.09 ± 0.7 73.00 ± 0.6 84.97 ± 0.6 73.08 ± 0.3 85.14 ± 4.9 84.27 ± 6.0 87.69 ± 3.3 37.69 ± 1.2 75.74
H2GCN (Zhu et al., 2020b) 78.12 ± 0.4 71.80 ± 0.8 84.26 ± 0.6 70.97 ± 0.2 84.86 ± 7.2 82.70 ± 5.2 87.65 ± 4.9 35.70 ± 1.0 74.51
+HalfHop 79.23 ± 0.8 71.90 ± 0.9 82.65 ± 0.8 71.26 ± 0.3 85.02 ± 7.0 83.54 ± 5.9 86.22 ± 5.2 34.83 ± 1.2 74.33
+UniGAP 80.26 ± 0.6 73.10 ± 0.5 85.13 ± 0.6 73.19 ± 0.2 86.14 ± 6.5 84.96 ± 5.0 87.73 ± 4.8 36.42 ± 0.8 75.87
@@
+HalfHop 77.95 ± 1.0 72.30 ± 0.7 83.49 ± 0.8 71.44 ± 0.5 83.87 ± 6.0 84.79 ± 7.6 85.76 ± 6.1 35.46 ± 1.0 74.38
+UniGAP 80.21 ± 0.9 72.90 ± 0.6 85.18 ± 0.4 71.98 ± 0.4 85.65 ± 5.4 84.37 ± 6.3 86.57 ± 5.0 36.51 ± 0.9 75.42
4.2.2. The Degree of Over-smoothing (RQ2)
@@
4.2.3. Interpretability in Mitigating Over-smoothing (RQ3)
To better analyze the reasons for mitigating over-smoothing, we examine the ratio of intermediate node insertions. This analysis provides valuable insights into the root causes of over-smoothing. As shown in Figure 4, we identify the locations where intermediate nodes are added. We observe that new nodes tend to be inserted between inter-class node pairs, slowing down the message passing between different classes and preventing the features of different classes from collapsing into identical vectors, thus mitigating the over-smoothing problem. For the portion of nodes inserted into intra-class edges, we speculate that some intra-class edges contain noise. Equipped with UniGAP, these harmful noises can be eliminated, thereby improving performance.
@@ 4.3. Ablation Study (RQ4)
In Section 3.3, we provide various strategies for instantiating our modules. In this subsection, we will explore the impact of different strategies within the various modules of UniGAP on model performance.
Comparison of Different Strategies in Trajectory Precomputation. In this study, we investigate three readily available strategies in Section 3.3.1: Zero Initialization, Non-parametric Message Passing, and Pretrained GNN. We conduct experiments on GCN+UniGAP across several datasets. From Figure 5, we observe that Pretrained GNN consistently outperforms both Zero Initialization and Non-parametric Message Passing. Due to its larger number of parameters, it can learn unique, additional information from the graph, allowing UniGAP to optimize the graph structure from a more advantageous starting point.
Figure 5. The performance of UniGAP with different Trajectory precomputation strategies (pink color), with different MVC Encoder (blue color), and with diiferent initialization for inserted nodes (green color).
@@
Comparsion of Different Initialization for Inserted Nodes. In Section 3.3.3, we introduce three strategies for initializing features for inserted nodes: (1) zero initialization, (2) mean feature of source and target node, and (3) learnable weighted sum of source and target node features. As shown in Figure 4, the learnable weighted sum achieves the best performance because it can adaptively adjust the weights of source and target features, thus generating better inserted node features.
@@ 4.4. Enhanced with LLMs (RQ5)
Recently, the graph domain has become closely connected with LLMs on text-attributed graphs (He et al., [n. d.]; Wang et al., 2024; Zhu et al., 2024a), demonstrating significant improvements due to the advanced capabilities in language understanding. In this subsection, we explore the potential of combining UniGAP with LLMs, like LLaMA2-7B (Touvron et al., 2023), on two text-attributed graphs, i.e., Arxiv and WikiCS (Mernyei and Cangea, 2020).
@@ 5. Conclusion
In this work, we propose UniGAP, a universal and adaptive graph upsampling methodology that alleviates the over-smoothing issue in the graph domain. By offering a universal framework that encompasses various graph upsampling techniques, UniGAP enhances the adaptability, performance, and interpretability of GNNs. Furthermore, we integrate UniGAP with LLMs, demonstrating the significant potential of combining LLMs with our method. For more intricate combinations, we leave it as future work. This research lays a robust foundation for future advancements in graph upsampling, offering valuable insights and practical solutions to longstanding challenges in the field.

### v2

@@ 5.2.1. Experiments on Small and Medium-scale Datasets
In summary, compared to all the baselines and datasets, UniGAP demonstrates a significant improvement, with an average absolute boost ranging from 1.4% to 6.7%, surpassing the SOTA baselines, showcasing our effectiveness.
@@
Homophily 0.66 0.63 0.77 0.42 0.00 0.02 0.05 0.01 Avg
GT (Dwivedi and Bresson, 2020) 79.08 ± 0.4 70.16 ± 0.8 82.22 ± 0.6 70.63 ± 0.4 84.18 ± 5.4 80.16 ± 5.1 82.74 ± 6.0 34.28 ± 0.7 72.92
GraphGPS (rampášek2023recipe) 79.94 ± 0.3 72.43 ± 0.6 82.44 ± 0.6 70.97 ± 0.4 82.21 ± 6.9 82.06 ± 5.1 85.36 ± 4.2 36.01 ± 0.9 73.92
NAGphormer (Chen et al., 2023) 80.57 ± 0.3 72.75 ± 0.8 84.20 ± 0.5 70.13 ± 0.6 80.12 ± 5.5 79.89 ± 7.1 82.97 ± 3.2 34.24 ± 0.9 73.11
@@
+DropEdge (Rong et al., 2020) 78.16 ± 0.6 72.12 ± 0.6 82.21 ± 0.5 71.87 ± 0.2 67.01 ± 4.0 58.74 ± 4.5 66.59 ± 6.1 30.72 ± 0.7 65.93
+AdaEdge (Chen et al., 2020a) 78.64 ± 0.7 72.15 ± 0.4 81.74 ± 0.4 71.98 ± 0.3 67.34 ± 4.2 62.08 ± 3.6 66.79 ± 6.5 30.87 ± 0.6 66.45
+HalfHop (Azabou et al., 2023) 77.42 ± 0.6 72.40 ± 0.7 81.36 ± 0.6 72.59 ± 0.2 71.62 ± 3.9 72.39 ± 4.8 66.14 ± 6.6 30.69 ± 0.7 68.07
+UniGAP (Ours) 80.06 ± 0.6 73.15 ± 0.6 84.20 ± 0.5 72.97 ± 0.2 77.37 ± 3.5 75.95 ± 4.7 69.59 ± 6.2 33.73 ± 0.9 70.87
GAT (Velickovic et al., 2018) 77.81 ± 0.4 70.89 ± 0.8 83.18 ± 0.5 71.95 ± 0.4 60.46 ± 6.2 58.22 ± 3.7 63.59 ± 6.1 30.36 ± 0.9 64.57
+DropEdge 77.15 ± 0.7 71.04 ± 0.9 83.20 ± 0.7 71.92 ± 0.5 69.45 ± 5.1 62.41 ± 5.6 68.47 ± 6.6 31.05 ± 0.8 66.84
+AdaEdge 75.24 ± 0.6 70.92 ± 0.9 82.67 ± 0.7 72.08 ± 0.2 75.04 ± 7.4 69.23 ± 5.3 71.52 ± 6.7 30.49 ± 0.9 68.40
+HalfHop 78.03 ± 0.5 71.83 ± 0.8 82.29 ± 0.8 72.56 ± 0.4 74.54 ± 6.1 66.48 ± 5.5 70.17 ± 6.0 30.28 ± 1.0 68.27
+UniGAP (Ours) 80.10 ± 0.3 73.20 ± 0.5 84.23 ± 0.6 74.06 ± 0.3 78.64 ± 4.9 76.04 ± 5.1 74.81 ± 5.3 32.79 ± 0.8 71.73
GraphSAGE (Hamilton et al., 2017) 78.67 ± 0.4 71.85 ± 0.6 83.76 ± 0.5 71.49 ± 0.3 82.43 ± 6.1 75.95 ± 5.3 81.18 ± 5.5 34.23 ± 1.0 72.44
+DropEdge 78.89 ± 0.7 71.72 ± 0.5 83.58 ± 0.6 71.42 ± 0.5 83.94 ± 5.7 77.42 ± 6.1 82.44 ± 4.9 34.75 ± 0.9 73.02
+AdaEdge 79.02 ± 0.7 71.07 ± 0.7 83.69 ± 0.6 70.24 ± 0.6 83.42 ± 5.9 79.81 ± 7.2 84.69 ± 5.4 34.97 ± 0.8 73.36
+HalfHop 78.94 ± 0.3 72.07 ± 0.8 84.27 ± 0.6 72.93 ± 0.5 85.95 ± 6.4 74.60 ± 6.0 85.88 ± 4.0 36.82 ± 0.7 73.92
+UniGAP (Ours) 80.21 ± 0.7 73.60 ± 0.3 86.46 ± 0.4 73.71 ± 0.2 86.52 ± 4.8 83.21 ± 5.7 86.98 ± 3.2 37.14 ± 0.8 75.98
@@
+DropEdge 80.18 ± 0.9 72.49 ± 0.7 84.74 ± 0.5 72.80 ± 0.7 79.55 ± 6.2 79.23 ± 6.5 80.01 ± 4.7 35.85 ± 1.0 73.11
+AdaEdge 79.62 ± 0.7 72.60 ± 0.6 85.07 ± 0.7 70.62 ± 0.4 78.09 ± 7.2 76.54 ± 4.8 80.47 ± 4.0 36.11 ± 0.9 72.39
+HalfHop 78.96 ± 0.9 72.43 ± 0.7 84.91 ± 0.6 72.91 ± 0.3 79.84 ± 6.1 80.04 ± 5.1 82.96 ± 3.7 36.01 ± 0.9 73.50
+UniGAP (Ours) 80.79 ± 0.7 73.70 ± 0.4 86.22 ± 0.4 73.32 ± 0.3 82.08 ± 4.2 81.16 ± 5.0 84.06 ± 3.8 36.47 ± 1.1 74.76
@@
+UniGAP (Ours) 80.09 ± 0.7 73.15 ± 0.6 84.97 ± 0.6 73.08 ± 0.3 85.14 ± 4.9 84.27 ± 6.0 87.69 ± 3.3 37.69 ± 1.2 75.74
H2GCN (Zhu et al., 2020a) 78.12 ± 0.4 71.80 ± 0.8 84.26 ± 0.6 70.97 ± 0.2 84.86 ± 7.2 82.70 ± 5.2 87.65 ± 4.9 35.70 ± 1.0 74.51
+DropEdge 78.32 ± 0.8 71.95 ± 0.9 83.17 ± 0.7 71.02 ± 0.6 85.11 ± 7.1 82.92 ± 5.1 87.91 ± 4.7 35.71 ± 1.2 74.51
+AdaEdge 78.29 ± 0.7 71.65 ± 0.7 83.31 ± 0.6 70.28 ± 0.5 83.79 ± 7.4 83.02 ± 5.0 87.38 ± 4.7 35.77 ± 1.1 74.19
+HalfHop 79.23 ± 0.8 71.89 ± 0.9 82.65 ± 0.8 71.26 ± 0.3 85.02 ± 7.0 83.54 ± 5.9 86.22 ± 5.2 34.83 ± 1.2 74.33
+UniGAP (Ours) 80.26 ± 0.6 73.10 ± 0.5 85.13 ± 0.6 73.19 ± 0.2 86.14 ± 6.5 84.96 ± 5.0 87.73 ± 4.8 36.42 ± 0.8 75.87
@@
+DropEdge 76.85 ± 1.1 71.74 ± 0.8 83.07 ± 0.7 71.09 ± 0.4 82.13 ± 6.7 82.92 ± 7.5 84.88 ± 6.0 35.21 ± 1.0 73.49
+AdaEdge 78.03 ± 1.1 70.82 ± 0.7 81.76 ± 0.8 70.99 ± 0.3 81.97 ± 6.2 84.32 ± 7.2 83.64 ± 6.6 35.06 ± 0.7 73.32
+HalfHop 77.95 ± 1.0 72.37 ± 0.7 83.49 ± 0.8 71.44 ± 0.5 83.87 ± 6.0 84.79 ± 7.6 85.76 ± 6.1 35.46 ± 1.0 74.38
+UniGAP (Ours) 80.21 ± 0.9 72.90 ± 0.6 85.18 ± 0.4 71.98 ± 0.4 85.65 ± 5.4 84.37 ± 6.3 86.57 ± 5.0 36.51 ± 0.9 75.42
5.2.2. Experiments on Large-scale Heterophilic Datasets
To evaluate the performance of UniGAP on large heterophilic datasets, we conducted experiments on Squirrel, Amazon-ratings, and Questions. As shown in Table 2, comparing with 2 scalable GNNs and NAGphormer, UniGAP achieves the best performance on all datasets, surpasses the runner-up over 1% absolute improvement on Amazon2M, demonstrating its effectiveness for large-scale graphs.
This further validates the effectiveness of decoupling multi-view attention and message interaction in capturing information and learning implicit long-range dependencies at a large scale among numerous nodes.
Table 2. UniGAP maintains its superiority on large heterophilic datasets, demonstrating strong generalization capabilities.
Dataset Squirrel Amazon-ratings Questions
Homophily 0.03 0.14 0.02
Metric Accuracy Accuracy ROC-AUC
GCN 45.87±2.3 53.80±0.6 79.02±0.6
   + AdaEdge 54.79±1.8 56.92±1.1 80.85±0.8
   + Halfhop 60.24±2.4 55.19±1.6 79.60±1.1
   + UniGAP(Ours) 65.58±1.6 59.76±0.8 82.73±0.9
GraphSAGE 43.78±1.9 55.40±0.2 77.21±1.3
   + AdaEdge 55.29±1.2 58.84±1.3 80.17±1.0
   + Halfhop 54.83±2.7 58.97±1.5 78.62±1.7
   + UniGAP(Ours) 65.32±1.4 62.83±0.9 84.21±0.8
H2GCN 43.62±2.3 56.81±0.9 79.28±0.8
   + AdaEdge 54.79±1.8 58.77±1.2 81.34±0.7
   + Halfhop 60.24±2.4 59.61±1.0 78.96±1.9
   + UniGAP(Ours) 69.67±1.6 63.75±1.1 83.59±0.7
5.3. The Degree of Over-smoothing (RQ2)
@@
Moreover, we observe that adding UniGAP on the baseline GCN substantially slows down the decay rate of MAD, roughly to about half of the original rate. For example, with UniGAP, the MAD value of at the 6 layer is comparable to the smoothness level of the baseline GCN at the 3 layer, which is consistent with the theoretical analysis in Section 4.2. On heterophilic datasets such as Texas and Cornell, the effect of UniGAP is more pronounced than on the homophilic Cora dataset, because heterophilic graphs are more susceptible to over-smoothing and therefore benefit more from slower propagation and a better adapted graph structure.
5.4. Interpretability in Mitigating Over-smoothing (RQ3)
To better analyze the reasons for mitigating over-smoothing, we examine the ratio of intermediate node insertions on a wide range of datasets. This analysis provides valuable insights into the root causes of over-smoothing. As shown in Figure 4, when UniGAP learned the augmented graph that achieved optimal performance, we identify the locations where intermediate nodes are added. We observe that new nodes tend to be inserted between inter-class node pairs, especially on heterophilic graphs where inter-class edges are more prevalent. This behavior slows down message passing across class boundaries and helps prevent the features of different classes from collapsing into similar vectors, thereby mitigating over-smoothing.
We also observe that a non-negligible fraction of new nodes is inserted on intra-class edges. One hypothesis is that some of these intra-class edges may in fact be noisy or weakly informative connections, and that inserting intermediate nodes on such edges effectively attenuates their influence and can lead to performance gains. A more systematic investigation of this hypothesis is left as a direction for future work. Overall, UniGAP highlights key structural bottlenecks where over-smoothing tends to occur, offering an interpretable view of how adaptive upsampling reshapes the graph to alleviate this phenomenon.
@@ 5.5. Ablation Study (RQ4)
In Section 3.1, 3.2 and 3.3, we provide various strategies for instantiating our modules. In this subsection, we will explore the impact of different strategies within the various modules of UniGAP on model performance.
Comparison of Different Strategies in Trajectory Precomputation. In this study, we investigate three readily available strategies in Section 3.1: Zero Initialization, Non-parametric Message Passing, and Pretrained GNN. We conduct experiments on GCN+UniGAP across several datasets. From Figure 5, we observe that Pretrained GNN consistently outperforms both Zero Initialization and Non-parametric Message Passing. Due to its larger number of parameters, it can learn unique, additional information from the graph, allowing UniGAP to optimize the graph structure from a more advantageous starting point.
Moreover, we conduct a sensitivity analysis on the pretrained GNN, where we adopt a weaker pretraining setup. We quantify the quality of the pretrained encoder by its validation performance on the pretext task. As shown in Table 3, better pretraining consistently improves the downstream accuracy of UniGAP. However, even with very weak or no pretraining, UniGAP still outperforms purely heuristic baselines, suggesting that UniGAP is not overly fragile to the quality of the initial trajectories.
Table 3. Sensitivity of UniGAP to pretraining quality on the Texas dataset. The pretrain score is the validation performance on the pretext task (higher is better).
Pretraining setup Pretrain score UniGAP+GCN Accuracy
None (Zero init) 73.12±3.8
Weak pretraining (10 ep.) 0.61 75.28±2.9
Moderate pretraining (50 ep.) 0.72 75.66±2.7
Strong pretraining (100 ep.) 0.78 75.73±2.7
Best pretraining (130 ep.) 0.81 75.85±2.1
Heuristic baseline (AdaEdge) 62.08±3.6
Figure 5. The performance of UniGAP with different Trajectory precomputation strategies (pink color), with different MVC Encoder (blue color), and with different initialization for inserted nodes (green color).
@@
Moreover, to verify whether the MVC encoder is truly learning the propensity to over-smoothing from trajectories rather than simply extracting generic features that are useful for the downstream classification task, we design a diagnostic experiment. Specifically, we replace the trajectory inputs with only the initial node features and keep the rest of the MVC and UniGAP pipeline unchanged. As shown in Table 4, this modification leads to a clear performance degradation compared to using full trajectories, indicating that the MVC encoder is indeed leveraging information about the evolution of node representations across layers, rather than acting as a generic feature extractor.
Table 4. Diagnostic experiment on the role of trajectories in the MVC encoder. We compare using full layer-wise trajectories versus using only the initial node features as input to MVC, while keeping the rest of the GCN+UniGAP pipeline unchanged.
Dataset MVC Input Test Accuracy
Cora Trajectories 84.20±0.5
Initial features only 81.37±0.7
Texas Trajectories 77.37±3.5
Initial features only 71.52±3.7
Comparsion of Different Initialization for Inserted Nodes. In Section 3.3, we introduce three strategies for initializing features for inserted nodes: (1) zero initialization, (2) mean feature of source and target node, and (3) learnable weighted sum of source and target node features. As shown in Figure 4, the learnable weighted sum achieves the best performance. We hypothesize that this is because it encodes a natural inductive bias for node insertion: the feature of an intermediate node should lie between its two endpoints in the representation space, but the optimal position is not necessarily the arithmetic mean. By learning asymmetric weights for the two endpoints, the model can adaptively shift the inserted node closer to the more informative side while still preserving local semantic consistency and requiring only a small number of additional parameters. In contrast, zero initialization places the new node far from the data manifold, making it harder for the downstream GNN to integrate it meaningfully, and simple averaging ignores potential asymmetry between the connected nodes.
Beyond these simple choices, more expressive initialization strategies are also possible. For example, one could treat the features of inserted nodes as independent learnable embeddings. We leave a systematic exploration of these richer generative initializations for inserted nodes as an interesting direction for future work.
@@ 5.6. Enhanced with LLMs (RQ5)
Recently, the graph domain has become closely connected with LLMs on text-attributed graphs (He et al., ; Wang et al., 2024; Zhu et al., 2024a), demonstrating significant improvements due to the advanced capabilities in language understanding. In this subsection, we explore the potential of combining UniGAP with LLMs, like LLaMA2-7B (Touvron et al., 2023), on two text-attributed graphs, i.e., Arxiv and WikiCS (Mernyei and Cangea, 2020).
@@ 6. Conclusion
In this work, we propose UniGAP, a universal and adaptive graph upsampling methodology that alleviates the over-smoothing issue in the graph domain. UniGAP enhances the adaptability, performance, and interpretability of GNNs. Furthermore, we integrate UniGAP with LLMs, demonstrating the significant potential of combining LLMs with our method. Despite its notable performance and interpretability, there are areas where UniGAP could be limited or further refined. First, although the additional time and space complexity introduced by UniGAP is moderate, applying UniGAP to very large graphs remains challenging. In particular, storing the features of the newly inserted nodes incurs extra memory overhead, which can become substantial on massive graphs. Second, while UniGAP mitigates over-smoothing by slowing propagation through adaptive node insertion, the induced increase in path length may exacerbate over-squashing in graphs where long-range dependencies are critical. Understanding and balancing this trade-off in a more principled way is an important direction for future work. Third, in this work we focus on node classification for studying over-smoothing. However, the UniGAP framework can be extended to other tasks such as link-prediction and graph classification. Exploring these extensions is an interesting direction for future work. In conclusion, we envision UniGAP to inspire further exploration of graph upsampling methods.